Find the area under the curve and estimates for the mean of 200 with standard deviation...

  1. Find the area under the curve and estimates for the mean of 200 with standard deviation 30. Find the following and draw the required area.
  1. P(x<155)
  2. P(x<228)
  3. Between P(155<x<228)
  4. Px≥228
Answer #1

Answer a) 0.0668

Step 1: Sketch the curve. The probability that X < 155 is equal to the blue area under the curve. 0.016 0.014 0.012 0.010 0.0

Answer b) 0.8238

Step 1: Sketch the curve. The probability that X<228 is equal to the blue area under the curve. 0.016 0.014 0.012 0.010 0.008

Answer c) 0.7570

P(155<x<228) = P(x<228) - P(x<155) = 0.8238 - 0.0668

P(155<x<228) = 0.7570

Answer d) 0.1762

P(x228) = 1 - P(x<228) = 1 - 0.8238

P(x228) = 0.1762